20 ml of 0.25N of strong acid and 30ml of 0.2N of strong base are mixed; the resulting solution is
  1. 0.02N acidic
  2. 0.025N basic
  3. 0.02N basic
  4. 0.025N acidic       and why?

Grade:

2 Answers

Sher Mohammad IIT Delhi
9 years ago
acid equivalents of solution N1*V1=20*.25=5
basic eqivalents of the solution N2*V2=30*0.2=6
so solution is basic N2*V2-N1*V1=1
N=1/50=0.02N basic (C)

sher Mohammad
faculty askiitians
911
9 years ago
Why not acidic?